OpenStudy (anonymous):

how do u solve 2(n+5)=-2

OpenStudy (anonymous):

distribute the 2 first

OpenStudy (anonymous):

first you multiply what is inside the (.....) and then you solve for n

OpenStudy (nincompoop):

a(x+4) = y ax + 4a = y ax + 4a +(-4a) = y + (-4a) ax = y - 4a x = (y/a) - 4

OpenStudy (anonymous):

2(n+5)=-2 2n+10=-2 2n=-2-10 2n=-12 n=-12divided by 2 n=-6

OpenStudy (nincompoop):

please do not give out answers

OpenStudy (anonymous):

why

OpenStudy (anonymous):

yeah i know i am sorry

OpenStudy (nincompoop):

because that is not really helping the person understand the concept.

OpenStudy (anonymous):

can u help me in something plz

OpenStudy (anonymous):

I am on 9th iand i need a experiment not a demostration can someone give me some plz